# Now at ]Max-Planck-Institut für Gravitationsphysik (Albert-Einstein-Institut) and Leibniz Universität Hannover, 30167 Hannover, Germany Now at ]Max-Planck-Institut für Gravitationsphysik (Albert-Einstein-Institut) and Leibniz Universität Hannover, 30167 Hannover, Germany

### v1

@@ I.1 ALPS II Design
Figure 1: Side view of the design of the full ALPS II experimental system. The HPL beam and PC spatial eigenmode is shown in red, while the BSM field traveling to the right is shown in blue. The dashed purple lines show the projection of the RC spatial eigenmode. A HWP before the PC can be used to configure the polarization state of the HPL with respect to the direction of the magnetic field.
@@
The spectral component of the field overlap $\eta_{\hat{z}}$, also referred to as the longitudinal field overlap, is discussed in more detail Section IV.1.4. It is dependent on the resonance condition of the cavity and can be approximated by the cavity Lorentzian:
@@ I.2 Optical System
The phase stability of all of the interference beatnotes is critical to the ALPS II heterodyne detection scheme. Therefore, the oscillators used to demodulate the science and veto signals, as well as those used for the PLLs, are all synchronized to a global clock and left to run uninterrupted for the duration of a science run. To set the frequencies relative to the RC FSR (referred to from here on as $f_{0}$), $f_{0}$ is measured and used to fix the beatnote frequencies before the run. Changes to the length of the RC that push its FSR away from the initial value of $f_{0}$ can introduce phase offsets to the measurement. This is discussed in more detail in Section IV.1.4. Another point discussed in more detail in Appendix A.1, is that because the oscillators are driven by digital clocks, they cannot use any arbitrary frequency, but must instead use a frequency that can be represented by an integer fraction of the clock frequency.
@@ I.3 Heterodyne Sensing
ALPS II utilizes a heterodyne detection system in which the LO is used as the local oscillator laser and is interfered with the signal field exiting the RC at RC1. The power of the resulting field traveling away from the curved mirror of the cavity toward the science PD can be expressed by the following equation.
$P_{\!{}_{\rm out}}(t)=P_{\!{}_{\rm LO}}+P_{\!{}_{\rm S}}+2\eta_{\!{}_{xy,{\rm LO}}}\sqrt{P_{\!{}_{\rm LO}}P_{\!{}_{\rm S}}}\cos{\{2\pi(\nu_{\!{}_{\rm LO}}-\nu_{\!{}_{\rm S}})t+\theta(t)}\}+$... (12)
@@
Due to several diagnostic pick-off mirrors, as well as the imperfect reflectivity and aperture of the mirrors in the path to the cavity, there are some optical losses as the light travels from the RC to the science PD. The power transmissivity on this path $T_{\!{}_{\rm sci}}$, was measured to be between 80% and 90% throughout the first science campaign. The power incident on the photodetector results in a voltage output by the photodetector that is proportional to its conversion gain $G$. The static voltage is removed with an analog high-pass filter and the resulting voltage, approximated by
@@
$=\frac{V_{\!{}_{\rm AC}}}{2}\left(e^{i2\pi m\frac{f_{\!{}_{\rm S}}}{f_{{}_{m_{1}}}}+\theta[m]}+e^{-i2\pi m\frac{f_{\!{}_{\rm S}}}{f_{{}_{m_{1}}}}-\theta[m]}\right)\,e^{i2\pi(f_{\!{}_{\rm S}}+2.4\,{\rm Hz})\frac{m}{f_{{}_{m_{1}}}}}+$... (16)
@@ I.4 Central Optical Bench
The COB has two light-tight boxes that sub-divide it into three areas. The HPL area on the left, the RC area on the right, and the intermediate RL area in between them. The purpose of this design is to contain the stray light from th HPL as well as possible and try to prevent it from coupling to the RC, while also allowing the HPL to be phase locked to the AL. The light-tight boxes therefore cannot be completely sealed in order to allow for the stabilization of the HPL frequency as well as open shutter measurements. To accommodate this they are equipped with the highly reflective (HR) mirrors LT1 and LT2. Because there are also light-tight boxes outside the vacuum system connected to the exit flanges of the light-tight boxes inside the vacuum system, the only way in and out of them is through the magnet bore or LT1 and LT2. Each has a transmissivity of $12.4\pm 0.4$ ppm for $p$-polarized light with a wavelength of 1064 nm at room temperature using an angle of incidence of $35^{\circ}$. The low transmissivity was chosen to help suppress the stray light signals that pass between the areas of the experiment.
@@
When the shutter is open, the HPL must pass through mirrors LT1, MZ3, MZ4 ($T_{\!{}_{\rm MZ4}}=275\pm 10$ ppm), LT2 before being incident on the RC. The combined transmissivity of these optics, and hence $T_{\!{}_{\rm COB}}$, was measured to be $T_{\!{}_{\rm COB}}=1.9\pm 0.1\times 10^{-17}$. These measurements are discussed in more detail in Section 5. With the shutter open, the HPL incident on the RC was roughly $6\times 10^{-16}$ W when 30 W of HPL power was used. Including mirror RC2, the on-axis optical attenuation between the two magnet strings was $T_{\!{}_{\rm COB}}\cdot T_{\!{}_{\rm RC2}}=9.7\pm 1.2\times 10^{-23}$ when the shutter was open. When considering the transmissivity of the RC and the coupling of the HPL to its eigenmodes, on the order of 10 aW of HPL power is expected at the science detector when the shutter is open.
@@ II Closed-Shutter Calibration
The open and closed shutter powers cannot be measured simultaneously. However, because the misalignment of the system due to drifts typically only became significant over several days, it was possible to calibrate the closed shutter data using the nearest neighboring open shutter periods. Open shutter measurements were performed daily during $\rm S_{{}_{\perp}}$ and several times per day for $\rm S_{{}_{\parallel}}$, to minimize the drifts between open and closed shutter periods. For $\rm S_{{}_{\perp}}$ this corresponded to 19 open shutter periods during the 15 day run, while for $\rm S_{{}_{\parallel}}$ 78 open shutter periods were performed over 41 days. The average time between the closed shutter data and its nearest open shutter period was roughly 8 hours for both runs. The calibration procedure is explained in further detail in Section II.4.
It should be noted that the purpose of the calibration procedure is to provide the results in terms of the electromagnetic field to BSM field conversion rate $\mathcal{P}_{\gamma\leftrightarrow\phi}$ and therefore this manuscript does not consider the effect of the magnet strings on the coupling of the BSM fields. Exclusion limits derived from these results (presented in Section III.4) on the coupling between electromagnetic fields and pseudoscalar, scalar, vector, and tensor bosons can instead be found in Ref. [7].
@@ II.3 Normalization by HPL Power
$P_{\phi}=\eta^{2}\beta_{\!{}_{\rm RC}}P_{i}\,\mathcal{P}_{\gamma\leftrightarrow\phi}^{2}$ (26)
@@
(a) $\rm S_{{}_{\perp}}$
(b) $\rm S_{{}_{\parallel}}$
@@ II.4 Closed to Open Shutter Ratio
In Section I.3, the integration of the time series of the $H$-function to produce the ‘$Z$-function’ was discussed. In order to remove systematic drifts and uncertainty from the coherent sum, the $H$-function must first be expressed in terms of the closed/open shutter ration before the closed shutter data series is integrated, by calculating the $Z$-function for each open shutter section of data. Then each closed shutter section of data is divided by its nearest valid point in the set of the open shutter data denoted by $C_{{\rm{open,}}j}$, which can be expressed as
@@
The magnitude squared and the phase of the open shutter measurements of both science and veto detector are shown in Figures 8(a), 8(b), 9(a), and 9(b), for $\rm S_{{}_{\perp}}$ and $\rm S_{{}_{\parallel}}$. Each data point represents the 19 and 57 distinct open shutter periods in $\rm S_{{}_{\perp}}$ and $\rm S_{{}_{\parallel}}$ respectively. These results show that $C_{{\rm{open,}}j}$ was extremely stable during $\rm S_{{}_{\perp}}$, especially the science detector, where the changes in $|C_{{\rm{open,}}j}|^{2}$ were less than 30% through the duration of the run and phase changes of roughly a tenth of a cycle. $C_{{\rm{open,}}j}$ appears to be less stable during $\rm S_{{}_{\parallel}}$. [^15] This is believed to be due to changes in the weather over the duration of the run. Here the main driver behind the changes in the magnitude and phase of $C_{{\rm{open,}}j}$ during $\rm S_{{}_{\parallel}}$ were the length changes of the RC detuning the HPL from the cavity resonance. For more details please refer to Section IV.1.4.
@@ III Results and Backgrounds
The results in terms of $|Z_{\rm closed}|^{2}$ are shown in Table II.6 with a measured value at the science detector of $(1.95\pm 0.26)\times 10^{-5}$ for $\rm S_{{}_{\perp}}$ and $(2.87\pm 0.36)\times 10^{-5}$ for $\rm S_{{}_{\parallel}}$. For the veto detector $|Z_{\rm closed}|^{2}$ was $(4.58\pm 0.53)\times 10^{-5}$ for $\rm S_{{}_{\perp}}$ and $(4.90\pm 0.57)\times 10^{-5}$ for $\rm S_{{}_{\parallel}}$. The error bars shown in Table II.6 represent the estimated systematic uncertainty that was calculated with the methods discussed in the previous section.
The resulting value of $|Z_{\rm closed}|^{2}$ for demodulation at the heterodyne frequencies given in Section I.3 lies in what we will refer to as the signal frequency bin, but the $Z$-function was also evaluated at alternative frequencies where no BSM contribution was to be expected. This allowed the backgrounds to be estimated without requiring background measurements with the magnets off, under the assumption that the neighboring bins are indicative of the background in the signal bin. The background estimation is summarized and compared with the measured data in Section III.2.
@@
Figure 10: $|Z(f)_{\rm closed}|^{2}$ measured by the science (a) veto (b) detectors during $\rm S_{{}_{\perp}}$ are shown in blue over the frequency range from 5-5 mHz to +5+5 mHz around the signal frequency bin. The mean technical noise is shown in green, while the shot noise limit is shown in black.
The results in the signal bin given in Table II.6, are consistent with the estimated background levels during both runs. Thus, neither run showed an excess statistically significant enough to claim the presence of a signal generated by new physics. Section III.3.1 therefore explains how the results and the background estimates are used to set exclusion limits on the conversion rate between electromagnetic and BSM fields.
@@ III.1 Alternative Frequency Analysis
Figure 11: $|Z(f)_{\rm closed}|^{2}$ measured by the science (a) veto (b) detectors during $\rm S_{{}_{\parallel}}$ are shown in blue over the frequency range from -5 mHz to +5 mHz around the signal frequency bin. The mean technical noise is shown in green, while the shot noise limit is shown in black.
@@ III.3.1 Detection Thresholds
(a) Histogram of the $W$ data from $\rm S_{{}_{\perp}}$
(b) Histogram of the $W$ data from $\rm S_{{}_{\parallel}}$
@@ III.4 Summary of Results
Table 4 gives a summary of the results $\rm S_{{}_{\perp}}$ and $\rm S_{{}_{\parallel}}$ in terms of $\mathcal{P}_{\gamma\leftrightarrow\phi}$. All results in this table are obtain using the Savitzky-Golay filter as this method is considered to be more robust than when the top-hat filter is used. Here the signal bin showed an electromagnetic power to BSM power conversion rate of $(4.4\pm 0.4)\times 10^{-14}$ for $\rm S_{{}_{\perp}}$ and $(5.3\pm 0.5)\times 10^{-14}$ for $\rm S_{{}_{\parallel}}$. The average estimated background at the detection system for $\rm S_{{}_{\perp}}$ was $(4.7\pm 0.4)\times 10^{-14}$ and for $\rm S_{{}_{\parallel}}$ was $(3.2\pm 0.3)\times 10^{-14}$. The estimated backgrounds were then used to calculate a $5\sigma$ detection threshold of $18\pm 2\times 10^{-14}$ for $\rm S_{{}_{\perp}}$ and $12\pm 1\times 10^{-14}$ for $\rm S_{{}_{\parallel}}$. In $\rm S_{\perp}$ no excess was seen above the estimated background. Although a slight excess with a significance of $1.8\,\sigma$ was observed in $\rm S_{\parallel}$ (the significance of the result for $\rm S_{\perp}$ was $0.8\,\sigma$), this is well below the $5\sigma$ detection threshold. Therefore, the discovery of new physics cannot be claimed for either measurement run. Signals with a $\mathcal{P}_{\gamma\leftrightarrow\phi}$ above $25\times 10^{-14}$ for $\rm S_{{}_{\perp}}$ and $17\times 10^{-14}$ for $\rm S_{{}_{\parallel}}$ can be excluded with a confidence level of 95%. These limits were calculated using the upper limit of the systematic uncertainty interval such that for $\rm S_{{}_{\perp}}$ $(22\pm 3)\times 10^{-14}\rightarrow 25\times 10^{-14}$ and for $\rm S_{{}_{\parallel}}$ $(15\pm 2)\times 10^{-14}\rightarrow 17\times 10^{-14}$.

### v2

# Any Light Particle Searches with ALPS II: Description of the first science campaign

@@ I.1 ALPS II Design
Figure 1: Side view of the design of the full ALPS II experimental system. The HPL beam and PC spatial eigenmode is shown in red, while the BSM field traveling to the right is shown in blue. The dashed purple lines show the projection of the RC spatial eigenmode. A HWP before the PC can be used to configure the polarization state of the HPL with respect to the direction of the magnetic field (shown as the black arrow).
@@
The spectral component of the field overlap $\eta_{\hat{z}}$, also referred to as the longitudinal field overlap, is discussed in more detail Section IV.2.1. It is dependent on the resonance condition of the cavity and can be approximated by the cavity Lorentzian:
@@ I.2 Optical System
The phase stability of all of the interference beatnotes is critical to the ALPS II heterodyne detection scheme. Therefore, the oscillators used to demodulate the science and veto signals, as well as those used for the PLLs, are all synchronized to a global clock and left to run uninterrupted for the duration of a science run. To set the frequencies relative to the RC FSR (referred to from here on as $f_{0}$), $f_{0}$ is measured and used to fix the beatnote frequencies before the run. Changes to the length of the RC that push its FSR away from the initial value of $f_{0}$ can introduce phase offsets to the measurement. This is discussed in more detail in Section IV.2.1. Another point discussed in more detail in Appendix A.1, is that because the oscillators are driven by digital clocks, they cannot use any arbitrary frequency, but must instead use a frequency that can be represented by an integer fraction of the clock frequency.
@@ I.3 Heterodyne Sensing
ALPS II utilizes a heterodyne detection system in which the LO is used as the local oscillator laser and is interfered with the signal field exiting the RC at RC1. The power of the resulting field traveling away from the curved mirror of the cavity toward the science PD can be expressed by
[…]
@@
Due to several diagnostic pick-off mirrors, as well as the imperfect reflectivity and aperture of the mirrors in the path to the cavity, there are some optical losses as the light travels from the RC to the science PD. The power transmissivity on this path $T_{\!{}_{\rm sci}}$, was measured to be between 80% and 90% throughout the first science campaign. The power incident on the photodetector results in a voltage output by the photodetector that is proportional to its conversion gain $G$. The static voltage is removed with an analog high-pass filter, suppressing the terms $P_{\!{}_{\rm LO}}$ and $P_{\!{}_{\rm S}}$ in Equation 12, and the resulting voltage, approximated by
@@
$=\frac{V_{\!{}_{\rm AC}}}{2}\left(e^{i\left(2\pi m\frac{f_{\!{}_{\rm S}}}{f_{{}_{m_{1}}}}+\theta[m]\right)}+e^{-i\left(2\pi m\frac{f_{\!{}_{\rm S}}}{f_{{}_{m_{1}}}}-\theta[m]\right)}\right)\,e^{i2\pi(f_{\!{}_{\rm S}}+2.4\,{\rm Hz})\frac{m}{f_{{}_{m_{1}}}}}+$... (16)
@@ I.4 Central Optical Bench
The COB has two light-tight boxes that sub-divide it into three areas. The HPL area on the left, the RC area on the right, and the intermediate RL area in between them. The purpose of this design is to contain the stray light from th HPL as well as possible and try to prevent it from coupling to the RC, while also allowing the HPL to be phase locked to the AL. The light-tight boxes therefore cannot be completely sealed in order to allow for the stabilization of the HPL frequency as well as open shutter measurements. To accommodate this they are equipped with the highly reflective (HR) mirrors LT1 and LT2. Because there are also light-tight boxes outside the vacuum system connected to the exit flanges of the light-tight boxes inside the vacuum system, the only way in and out of them is through the magnet bore or LT1 and LT2. Each has a transmissivity of $12.4\pm 0.4$ ppm for $p$-polarized ($\hat{x}$ direction) light with a wavelength of 1064 nm at room temperature using an angle of incidence of $35^{\circ}$. The low transmissivity was chosen to help suppress the stray light signals that pass between the areas of the experiment.
@@
When the shutter is open, the HPL must pass through mirrors LT1, MZ3, MZ4 ($T_{\!{}_{\rm MZ4}}=275\pm 10$ ppm), LT2 before being incident on the RC. The combined transmissivity of these optics, and hence $T_{\!{}_{\rm COB}}$, was measured to be $T_{\!{}_{\rm COB}}=1.9\pm 0.1\times 10^{-17}$. These measurements are discussed in more detail in Section 5. With the shutter open, the HPL incident on the RC was roughly $6\times 10^{-16}$ W when 30 W of HPL power was used. Including mirror RC2, the on-axis optical attenuation between the two magnet strings was $T_{\!{}_{\rm COB}}\cdot T_{\!{}_{\rm RC2}}=9.7\pm 1.2\times 10^{-23}$ when the shutter was open. When considering the transmissivity of the RC and the coupling of the HPL to its eigenmode, on the order of 10 aW of HPL power is expected at the science detector when the shutter is open.
@@ II Closed-Shutter Calibration
In using the ratio $P_{\phi}/P_{\rm open}$ to calibrate the results, we rely on the assumption that the COB optics will only attenuate the power of the HPL and not perturb the spatial eigenmode with respect to the eigenmode of the BSM field. There are a number of systematic effects that could invalidate this assumption and, as Section II.1 explores, several checks were in place specifically to ensure that this was not the case.
The open and closed shutter powers cannot be measured simultaneously. However, because the misalignment of the system due to drifts typically only became significant over several days, it was possible to calibrate the closed shutter data using the nearest neighboring open shutter periods. Open shutter measurements were performed daily during $\rm S_{{}_{\perp}}$ and several times per day for $\rm S_{{}_{\parallel}}$, to minimize the drifts between open and closed shutter periods. For $\rm S_{{}_{\perp}}$ this corresponded to 19 open shutter periods during the 15 day run, while for $\rm S_{{}_{\parallel}}$ 78 open shutter periods were performed over 41 days. The average time between the closed shutter data and its nearest open shutter period was roughly 8 hours for both runs. A critical component of this is ensuring that the system was in a stable state during the periods in which data was acquired for both the times when the shutter was open or closed. The process of identifying these times and flagging the data based on the status of the system is explained in Section II.2.
Because it was possible to continuously measure $P_{{}_{\rm HPL}}$, the closed and open shutter data was also normalized by the HPL power to further reduce the systematic uncertainty of the measurement, and this is described in Section II.3. The rest of the calibration procedure to calculate the result from the heterodyne detection system in terms of $Z$ and then calibrate this in $\mathcal{P}_{\gamma\leftrightarrow\phi}$ is explained in Section II.4. A second PD called the veto detector, was implemented on the opposite side of the RC from the science detector, to verify potential signals and we examine this in Section II.5. The uncertainty in the calibration of the ratio $P_{\phi}/P_{\rm open}$ and the results in terms of $\mathcal{P}_{\gamma\leftrightarrow\phi}$ is then scrutinized in Sections II.6 and II.7 respectively.
It should be noted that the purpose of the calibration procedure is to provide the results in terms of the electromagnetic field to BSM field conversion rate $\mathcal{P}_{\gamma\leftrightarrow\phi}$ and therefore this manuscript does not consider the effect of the magnet strings on the coupling of the BSM fields. Exclusion limits, derived from the results presented in Section III.4, on the coupling between electromagnetic fields and pseudoscalar, scalar, vector, and tensor bosons can instead be found in Ref. [7].
@@ II.3 Normalization by HPL Power
$P_{\phi}=\eta^{2}\beta_{\!{}_{\rm RC}}P_{\rm i}\,\mathcal{P}_{\gamma\leftrightarrow\phi}^{2}$ (26)
@@
(a) HPL power ($\rm S_{{}_{\perp}}$)
(b) HPL power ($\rm S_{{}_{\parallel}}$)
@@ II.4 Closed to Open Shutter Ratio
In Section I.3, the integration of the time series of the $H$-function to produce the ‘$Z$-function’ was discussed. In order to remove systematic drifts and uncertainty from the coherent sum, the $H$-function must first be expressed in terms of the closed/open shutter ratio before the closed shutter data series is integrated, by calculating the $Z$-function for each open shutter section of data. Then each closed shutter section of data is divided by its nearest valid point in the set of the open shutter data denoted by $C_{{\rm{open,}}j}$, which can be expressed as
@@
The magnitude squared and the phase of the open shutter measurements of both science and veto detector are shown in Figures 8(a), 8(b), 9(a), and 9(b), for $\rm S_{{}_{\perp}}$ and $\rm S_{{}_{\parallel}}$. Each data point represents the 19 and 57 distinct open shutter periods in $\rm S_{{}_{\perp}}$ and $\rm S_{{}_{\parallel}}$ respectively. These results show that $C_{{\rm{open,}}j}$ was extremely stable during $\rm S_{{}_{\perp}}$, especially the science detector, where the changes in $|C_{{\rm{open,}}j}|^{2}$ were less than 30% through the duration of the run and phase changes of roughly a tenth of a cycle. $C_{{\rm{open,}}j}$ appears to be less stable during $\rm S_{{}_{\parallel}}$. [^15] This is believed to be due to changes in the weather over the duration of the run. Here the main driver behind the changes in the magnitude and phase of $C_{{\rm{open,}}j}$ during $\rm S_{{}_{\parallel}}$ were the length changes of the RC detuning the HPL from the cavity resonance. For more details please refer to Section IV.2.1.
@@ III Results and Backgrounds
The results acquired by the heterodyne detection system using the science and veto detectors are shown for both runs in terms of $|Z_{\rm closed}|^{2}$ in Table II.6. The measured value given by the science detector data was $(1.95\pm 0.26)\times 10^{-5}$ for $\rm S_{{}_{\perp}}$ and $(2.87\pm 0.36)\times 10^{-5}$ for $\rm S_{{}_{\parallel}}$. For the veto detector $|Z_{\rm closed}|^{2}$ was $(4.58\pm 0.53)\times 10^{-5}$ for $\rm S_{{}_{\perp}}$ and $(4.90\pm 0.57)\times 10^{-5}$ for $\rm S_{{}_{\parallel}}$. The error bars shown in Table II.6 represent the estimated systematic uncertainty that was calculated with the methods discussed in the previous section.
The resulting values of $|Z_{\rm closed}|^{2}$ in Table II.6 lie in what we will refer to as the signal frequency bin, because the data was demodulated at the heterodyne frequencies given in Section I.3. As Section III.1 explains, the $Z$-function was also evaluated at alternative frequencies where no BSM contribution was expected. This allowed the backgrounds to be estimated without requiring background measurements with the magnets off, under the assumption that the neighboring bins are indicative of the background in the signal bin. The estimated background in the signal bin is shown as the second and third columns of Table II.6and this process is summarized and compared with the measured data in Section III.2.
@@
Figure 10: $|Z(f)_{\rm closed}|^{2}$ measured by the science (a) and veto (b) detectors during $\rm S_{{}_{\perp}}$ are shown in blue over the frequency range from 5-5 mHz to +5+5 mHz around the signal frequency bin. The mean technical noise is shown in green, while the shot noise limit is shown in black.
Section III.3 explains how the statistical distribution of the background was calculated by comparing the estimated background to the measured values at frequencies outside the signal bin. This distribution was then used to set detection thresholds that determined whether or not a signal was present at a given confidence level.
The results in the signal bin given in Table II.6 are consistent with the estimated background levels during both runs. Thus neither run showed an excess statistically significant enough to claim the presence of a signal generated by new physics. Section III.3 therefore also examines how exclusion limits on the conversion rate between electromagnetic and BSM fields were set based on the background distribution. The resulting exclusion limits are summarized in terms of $\mathcal{P}_{\gamma\leftrightarrow\phi}$ in Section III.4.
@@
Measurements of the transmissivities of the COB optics are then described in Section 5, as these parameters are critical to calculating $\mathcal{P}_{\gamma\leftrightarrow\phi}$ from the results of $|Z_{\rm closed}|^{2}$.
@@ III.1 Alternative Frequency Analysis
Figure 11: $|Z(f)_{\rm closed}|^{2}$ measured by the science (a) and veto (b) detectors during $\rm S_{{}_{\parallel}}$ are shown in blue over the frequency range from -5 mHz to +5 mHz around the signal frequency bin. The mean technical noise is shown in green, while the shot noise limit is shown in black.
@@ III.3.1 Detection Thresholds
(a) Histogram of the $W$ data ($\rm S_{{}_{\perp}}$)
(b) Histogram of the $W$ data ($\rm S_{{}_{\parallel}}$)
@@ III.4 Summary of Results
Table 4 gives a summary of the results $\rm S_{{}_{\perp}}$ and $\rm S_{{}_{\parallel}}$ in terms of $\mathcal{P}_{\gamma\leftrightarrow\phi}$. All results in this table are obtain using the Savitzky-Golay filter as this method is considered to be more robust than when the top-hat filter is used. Here the signal bin showed an electromagnetic power to BSM power conversion rate of $(4.4\pm 0.4)\times 10^{-14}$ for $\rm S_{{}_{\perp}}$ and $(5.3\pm 0.5)\times 10^{-14}$ for $\rm S_{{}_{\parallel}}$. The average estimated background at the detection system for $\rm S_{{}_{\perp}}$ was $(4.7\pm 0.4)\times 10^{-14}$ and for $\rm S_{{}_{\parallel}}$ was $(3.2\pm 0.3)\times 10^{-14}$. The estimated backgrounds were then used to calculate a $5\sigma$ detection threshold of $(18\pm 2)\times 10^{-14}$ for $\rm S_{{}_{\perp}}$ and $(12\pm 1)\times 10^{-14}$ for $\rm S_{{}_{\parallel}}$. In $\rm S_{\perp}$ no excess was seen above the estimated background. Although a slight excess with a significance of $1.8\,\sigma$ was observed in $\rm S_{\parallel}$ (the significance of the result for $\rm S_{\perp}$ was $0.8\,\sigma$), this is well below the $5\sigma$ detection threshold. Therefore, the discovery of new physics cannot be claimed for either measurement run. Signals with a $\mathcal{P}_{\gamma\leftrightarrow\phi}$ above $25\times 10^{-14}$ for $\rm S_{{}_{\perp}}$ and $17\times 10^{-14}$ for $\rm S_{{}_{\parallel}}$ can be excluded with a confidence level of 95%. These limits were calculated using the upper limit of the systematic uncertainty interval such that for $\rm S_{{}_{\perp}}$ $(22\pm 3)\times 10^{-14}\rightarrow 25\times 10^{-14}$ and for $\rm S_{{}_{\parallel}}$ $(15\pm 2)\times 10^{-14}\rightarrow 17\times 10^{-14}$.